OpenStudy (anonymous):

Solve the equation for y. 4x – 5y = 9

OpenStudy (shamil98):

solve the equation then. get y isolated.

jimthompson5910 (jim_thompson5910):

Subtract 4x from both sides (step 1). Then divide both sides by -5 (step 4) to fully isolate y \[\Large 4x - 5y = 9\] \[\Large -5y = 9-4x\] \[\Large -5y = -4x+9\] \[\Large y = \frac{-4x+9}{-5}\] \[\Large y = \frac{-4x}{-5}+\frac{9}{-5}\] \[\Large y = \frac{4}{5}x-\frac{9}{5}\]

OpenStudy (anonymous):

thanks jim c:

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

1. Add 4x to both sides ( -5y = 4x + 9) 2. Divide both sides by -5 ( y = -4/5x - 9/5) Whatever you do to one side of the equation you have to do to the other side to keep it balanced. Most algebra equations work like this.